OpenStudy (anonymous):

what is the maximum product of two numbers whose sum is -4?

OpenStudy (anonymous):

x+y =-4 x=-y - 4 = -(y+4) x y = -(y+4) y use x = 0 , y=-4 and get 0 as product, which is maximum not negative. Doesn't limit the absolute value of the product significantly, could be infinite

OpenStudy (anonymous):

If the two numbers are x and y, I think the answer is x = -2 and y = -2. Normally a square shape maximizes area. It's a common type of question. x + y = -4 We want to know for x times y, what is the maximum? x = (-y -4) from above. xy becomes (-y-4)(y)= -y^2 - 4y Do you know how to do derivatives yet?

OpenStudy (anonymous):

4 times 0 is 0 (-2)(-2) = 4

OpenStudy (anonymous):

I think @jam133 is right and I am wrong. Sorry.

OpenStudy (anonymous):

It is a tricky question!